OpenStudy (anonymous):

Can Someone Help? <3 3/4x + 5/8 = 4x

OpenStudy (mertsj):

\[\frac{3}{4}x+\frac{5}{8}=4x\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Multiply every term by 8

OpenStudy (anonymous):

what do u mean

OpenStudy (mertsj):

Do you see where it says 3/4 x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do i multiply 3x8 and 4x8

OpenStudy (mertsj):

\[\frac{3}{4}x \times8 +\frac{5}{8}\times8=4x \times 8\]

OpenStudy (anonymous):

im not good with fractions /,\

OpenStudy (mertsj):

That's why God created calculators.

OpenStudy (anonymous):

would i divide 3/4 then x 8 ?

OpenStudy (mertsj):